# Turbulent gas-rich discs at high redshift: origin of thick stellar discs through 3D ‘baryon sloshing’

### v1

@@ 6 Looking forward
In our first Nexus paper (Bland-Hawthorn et al., 2024), we showed how bulges, bars and spiral arms emerge even in turbulent, gas-rich environments, consistent with the remarkable new JWST results (Guo et al., 2022; Le Conte et al., 2023). In the present work, we focus on dynamical processes that unbind the disc and cause it to slosh around. The sloshing leads to a thickened stellar disc with properties remarkably consistent with our Galactic $\alpha$-enriched thick disc, in particular, the prediction of vertical gradients in age and metallicity, with only a weak to no radial gradient. These are complex non-linear processes that can only be addressed through numerical simulations presently. Analytic models that describe the evolution of fast and slow fluctuations are under way but at present they have been applied to only a few restricted cases (Hamilton et al., 2024; Binney, 2025).
There is little doubt that gas-dominated discs were abundant at early times. The most obvious manifestations of turbulent galaxies at high redshift are the observed high gas fractions and large gas clouds (e.g. Fensch & Bournaud, 2021), elevated star formation rates and broadened gas kinematics seen in ALMA and JWST surveys (Sec. 1). The enhanced star-forming discs at high redshift exhibit larger velocity dispersions than for local discs, regardless of the emission diagnostic used (Ejdetjärn et al., 2022), and presumably these reflect enhanced levels of turbulent energy.
@@
In a companion paper (Bland-Hawthorn et al., 2025), we explore both morphological and kinematic signatures of baryon sloshing. We highlight specific signatures in low-order ($m=1,2,3$) Fourier components and show that these effects may already be evident in high-redshift ALMA observations (e.g. Tsukui & Iguchi, 2021; Tsukui et al., 2024b). We focus our study on both the morphology and kinematics of the starlight and the cool gas for high-redshift galaxies up to $z\sim 4$. There are now dozens of galaxies that have been mapped by ALMA in cool gas out to this redshift, as discussed in Sec. 1. The notion of stellar kinematics beyond $z\sim 1$ may seem fanciful, except that at least two sources now have (sufficiently) resolved stellar velocity fields in the range $z=3-4$ (D’Eugenio et al., 2023; Pérez-González et al., 2024), with more to come. This was made possible by the remarkable sensitivity of the JWST NIRSpec integral field spectrograph. These are direct line-of-sight observations unaided by gravitational lensing. Reconstructed velocity fields from lensing are also now possible in the same redshift window. Consistent with our models, these stellar discs are relatively young ($300-700$ Myr) and thus detectable in Balmer absorption lines. For both discs, the star formation rates are of order 10 M yr-1 and considered ‘quiescent’ for such an early epoch, but these rates were an order of magnitude higher at even earlier times.
In future work, already under way, we will look at how merger activity affects our results during the gas-rich phase. It will be important to separate and quantify the respective roles of baryon sloshing and merger activity. We will evolve the Nexus models over the lifetime of the Milky Way. What are the likely chemodynamic signatures that survive? We conjecture that the different baryon components will have an excess of stars on retrograde orbits, induced by the sloshing, compared to other scenarios. While such orbits can arise from past mergers (e.g. Gallart et al., 2019; Amarante et al., 2020), there may be chemodynamic signatures that separate these two distinct processes.
@@ 7 Acknowledgments
JBH is indebted to Merton College and the Beecroft building for providing a stimulating research environment. We are particularly indebted to James Binney for an advanced copy of his new book Stellar Dynamics (2025), and for its innovative analysis of dynamical fluctuations that was pivotal for this work. Moreover, we received valuable feedback from seminars at Marseille, Paris, Shanghai, UCL, Cambridge and Oxford. Important questions were raised by Lia Athanassoula, Francoise Combes, Walter Dehnen, Wyn Evans and Vasily Belokurov, and we have sought to answer them here. We are grateful to Eugene Vasiliev for continued assistance and adaptions with agama.
@@ Appendix A Radial and kinematic signatures of baryon sloshing
Figure 15: Baryon sloshing presented as the change in radius $R_{\rm slosh}=R(x,y,z)$ with cosmic time (upper panels) in the reference frame of the total potential for $f_{\rm gas}=0,20,40,60,80,100$%. The lower panels are the derivative $\dot{R}$ of the upper panels. The red lines indicate the changing physical scales between panels, where the overall trend is increasing with $f_{\rm gas}$. Note also that the sloshing frequency declines with increasing $R$, as expected from Eq. 15.
Figure 16: (Left) Where new stars are born in time as a function of vertical height $|z|.$ The LHS panels look very similar for all $f_{\rm gas}$, quite unlike the RHS panels. (Right) The rms $z$-scale height for newborn stars as a function of radius $R$ (in kpc) for each value of $f_{\rm gas}$ for the halo accretion cases. In all plots, the colour coding indicates the time evolution, as indicated. The scale height of the $\alpha$-rich thick disc today at $R=8$ kpc are indicated by the blue star (Bland-Hawthorn & Gerhard, 2016). Note that the vertical scale has a fourfold stretch compared to the horizontal axis. The gas-rich models are able to reproduce the thickness of the $\alpha$-rich disc within $1.5-2$ Gyr.
Appendix B The case against thick discs from in situ star formation in a turbulent disc
In Sec. 5.1, we presented our evidence for disc heating being a gradual accumulation of energy from baryon sloshing within the dark matter potential. Star formation feedback and turbulence are implicated in the thickening of discs, but not from stars being born in situ far from the cold molecular disc. In Fig. 16, most stars in dynamical equilibrium are born close to the plane within the scale height of the cold gas. The remaining 20--30% of stars are injected under non-equilibrium conditions. The LHS figures are similar in their temporal behaviour for all $f_{\rm gas}$, unlike the RHS figures. In the RHS panel, the disc heating rate for newly formed stars in time increases rapidly with $f_{\rm gas}$. This is because the Brownian motion arising from baryon sloshing heats the stellar disc essentially isotropically, and the amplitude of this motion is a strong function of $f_{\rm gas}$.

### v2

@@ 6 Looking forward
In our first Nexus paper (Bland-Hawthorn et al., 2024), we showed how bulges, bars and spiral arms emerge even in turbulent, gas-rich environments, consistent with the remarkable new JWST results (Guo et al., 2023; Le Conte et al., 2024). In the present work, we focus on dynamical processes that unbind the disc and cause it to slosh around. The sloshing leads to a thickened stellar disc with properties remarkably consistent with our Galactic $\alpha$-enriched thick disc, in particular, the prediction of vertical gradients in age and metallicity, with only a weak to no radial gradient. These are complex non-linear processes that can only be addressed through numerical simulations presently. Analytic models that describe the evolution of fast and slow fluctuations are under way but at present they have been applied to only a few restricted cases (Hamilton et al., 2024; Binney, 2025).
There is little doubt that gas-dominated discs were abundant at early times. This has been evident for the past two decades, starting with early molecular gas observations and ground-based integral field studies of ionized gas (Chapman et al., 2004; Genzel et al., 2006; Förster Schreiber et al., 2006; Shapiro et al., 2008; Swinbank et al., 2012). The most obvious manifestations of turbulent galaxies at high redshift are the observed high gas fractions and large gas clouds (e.g. Fensch & Bournaud, 2021), elevated star formation rates and broadened gas kinematics seen in ALMA and JWST surveys (Sec. 1). The enhanced star-forming discs at high redshift exhibit larger velocity dispersions than for local discs, regardless of the emission diagnostic used (Ejdetjärn et al., 2022), and presumably these reflect enhanced levels of turbulent energy.
@@
In a companion paper (Bland-Hawthorn et al., 2025), we explore both morphological and kinematic signatures of baryon sloshing. We highlight specific signatures in low-order ($m=1,2,3$) Fourier components and show that these effects may already be evident in high-redshift ALMA observations (e.g. Tsukui & Iguchi, 2021; Tsukui et al., 2024b). We focus our study on both the morphology and kinematics of the starlight and the cool gas for high-redshift galaxies up to $z\sim 4$. There are now dozens of galaxies that have been mapped by ALMA in cool gas out to this redshift, as discussed in Sec. 1. The notion of stellar kinematics beyond $z\sim 1$ may seem fanciful, except that at least two sources now have (sufficiently) resolved stellar velocity fields in the range z=34z=3-4 (D’Eugenio et al., 2024; Pérez-González et al., 2025), with more to come. This was made possible by the remarkable sensitivity of the JWST NIRSpec integral field spectrograph. These are direct line-of-sight observations unaided by gravitational lensing. Reconstructed velocity fields from lensing are also now possible in the same redshift window. Consistent with our models, these stellar discs are relatively young (300700300-700 Myr) and thus detectable in Balmer absorption lines. For both discs, the star formation rates are of order 10 M yr-1 and considered ‘quiescent’ for such an early epoch, but these rates were an order of magnitude higher at even earlier times.
In future work, already under way, we will look at how merger activity affects our results during the gas-rich phase. It will be important to separate and quantify the respective roles of baryon sloshing and merger activity. We will evolve the Nexus models over the lifetime of the simulated Milky Way-like galaxies. What are the likely chemodynamic signatures that survive? We conjecture that the different baryon components will have different heating signatures, induced by the sloshing, compared to other scenarios. While such orbits can arise from past mergers (e.g. Gallart et al., 2019; Amarante et al., 2020), there may be chemodynamic signatures that separate these two distinct processes. If most galaxies are ultimately found to have chemically bimodal discs, given that most discs go through a gas-rich phase in the early Universe, the sloshing mechanism is arguably the most natural process to explain their prevalence.
@@ 7 Acknowledgments
JBH is indebted to Merton College and the Beecroft building for providing a stimulating research environment. We are particularly indebted to James Binney for an advanced copy of his new book Stellar Dynamics (2025), and for its innovative analysis of dynamical fluctuations that was pivotal for this work. We also thank him for providing the code for the calculation in Sec. 4.2. Moreover, we received valuable feedback from seminars delivered at Marseille, Paris, Shanghai, UCL, Cambridge and Oxford. Important questions were raised by Lia Athanassoula, Chris Hamilton, Francoise Combes, Bruce Elmegreen, Walter Dehnen, Wyn Evans and Vasily Belokurov, and we have sought to answer them here. We are grateful to Eugene Vasiliev and Romain Teyssier for continued assistance and adaptions with agama and Ramses respectively. We are also indebted to a thorough and perceptive referee.
@@ Appendix A Radial and kinematic signatures of baryon sloshing
Figure 16: Baryon sloshing presented as the change in radius $R_{\rm slosh}=\sqrt{x_{\rm slosh}^{2}+y_{\rm slosh}^{2}+z_{\rm slosh}^{2}}$ with cosmic time (upper panels)
in the reference frame of the total potential for $f_{\rm gas}=0,20,40,60,80,100$%. The lower panels are the derivative $\dot{R}_{\rm slosh}$ of the upper panels. The red lines indicate the changing physical scales between panels, where the overall trend is increasing with $f_{\rm gas}$. Note also that the sloshing frequency declines with increasing $R$, as expected from Eq. 15.